OpenStudy (anonymous):

The area of a square garden is 200 m². How long is the diagonal?

OpenStudy (anonymous):

show me how to do it don't just give the answer please

OpenStudy (hoblos):

area of a square = (side)^2 = 200 so the side of the square = √200 = √(100*2) = 10√2 diagonal= side*√2 = 10√2 * √2 = 10*2 = 20m

OpenStudy (anonymous):

thanks so much ! c:

OpenStudy (hoblos):

my pleasure xD